Topbaker Limited has just begun buying and selling Baking Powdeґ as an additional product to its main business of providing baking care goods and services. The powder is purchased in bulk and resold in smaller quantities to customers. Ben, the manager, supplies the following inventory information that relates to the month of January 2019. The company operates on an accruals basis and uses a perpetual, First in First Out (FIFO) inventory method (assume that all sales are on credit and ignore GST). Date Details (Baking Powder Inventory) 1-Jan Opening Balance: 30 kgs at $2.10/kg 5-Jan Purchases: 45 kgs at $2.20/kg 10-Jan Sales: 52 kgs at $2.80/kg 22-Jan Purchases: 34 kgs at $2.30/kg 31-Jan Sales: 19 kgs at $2.80 /kg Required: (a) Prepare an inventory card for the Baking Powder for the month ended 31 January 2019 (please disclose the total closing inventory balance in dollar terms) (b) Calculate the total cost of goods sold and gross profit from the sale of Baking Powder for the month ended 31 January 2019. Describe the effect on the gross profit and closing inventory balance for January 2019 if Topbaker Limited had used the weighted average cost method for recording inventory rather than the First In First Out method (no calculations required). (c)
Answer #1

(a)

Inventory card (For the month of January, 2019)

Purchases Sales Balance
Date Quantity Rate Value Quantity Rate Value Quantity Rate Value
Jan. 1 30 2.10 63
Jan. 5 45 2.20 99 30 2.10 63
45 2.20 99
Jan. 10 30 2.10 63
22 2.20 48.4 23 2.20 50.6
Jan. 22 34 2.30 78.2 23 2.20 50.6
34 2.30 78.2
Jan. 31 19 2.20 41.8 4 2.20 8.8
34 2.30 78.2

(b)

Cost of goods sold = 63 + 48.4 + 41.8

= $153.2

Sales = 52 x 2.80 + 19 x 2.80

= $198.8

Gross profit = Sales - Cost of goods sold

= 198.8 - 153.2

= $45.6

(c)

Inventory card (For the month of January, 2019) (Weighted average method)

Purchases Sales Balance
Date Quantity Rate Value Quantity Rate Value Quantity Rate Value
Jan. 1 30 2.10 63
Jan. 5 45 2.20 99 75 2.16 162
Jan. 10 52 2.16 112.32 23 2.16 49.68
Jan. 22 34 2.30 78.2 57 2.24 127.88
Jan. 31 19 2.24 42.56 38 2.24 85.32

Cost of goods sold = 112.32 + 42.56

= $154.88

Sales = 52 x 2.80 + 19 x 2.80

= $198.8

Gross profit = Sales - Cost of goods sold

= 198.8 - 154.88

= $43.92

If weighted average method is used, cost of goods sold will increase and gross profit will increase.

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.